A chemical storeroom has an 80% alcohol solution and a 30% alcohol solution. How many milliliters of each should be used to obtain 50 milliliters of a 60% solution.
Answer by rapaljer(4671)

Let x = amount of 80% solution
50-x= amount of 30% solution
50 = number of milliliters of 60%

.80x+.30(50-x)=.60*50
.80x+15-.30x=30
.50x+15=30
.50x=15
x=15/.50=30 milliliters of 80%
50-x=20 milliliters of 30%

Dr. Robert J. Rapalje